U.S five cent coins are made from a combination of two metals.Nickel and copper. For every pound of nickel, 3 lbs of copper are used. How many pounds of copper would be needed to make 560 lbs of five cent coins?
Answer by richwmiller(17219)

3/4*560=3/4*140*4=3*140=420